# Ranking Entities along Conceptual Space Dimensions with LLMs: An Analysis of Fine-Tuning Strategies

### v1

@@ 5 Experiments
Table 4 shows the 10 highest and lowest ranked entities, according to the rankings from the SVM method with the pairwise Llama2-13B model. The results for sweetness and saltiness were obtained with the model that was trained on WD+TG+Rocks. The rankings for scary and funny movies were obtained with the model that was trained on WD+Taste+Rocks. The ranking for population was obtained with the model that was trained on WD1. The table shows that the model was successful in selecting these top and bottom ranked entities. The top-ranked entities for sweetness, for instance, are all clearly sweet food items, while none of the bottom ranked entities are. Similar observations can be made for the other features. The model is sometimes less successful in distinguishing middle-ranked entities from bottom-ranked entities. For instance, most cheeses appear at the bottom of the ground truth ranking, whereas the model predicted these to be somewhere closer to the middle. [^10] A more detailed analysis of such errors can be found in the appendix. In the population example in Table 4, we can see that while the top-ranked entities are all countries with a high population, their relative ranking is not accurate. For instance, Nigeria is only the 6th most populous country in the dataset, but it is ranked in second place.
@@ 6 Conclusions
Acknowledgments
This work was supported by EPSRC grant EP/V025961/1.
@@ Appendix A Further Details
To train the three base models, i.e., meta-llama/Llama-2-7b-hf, mistralai/Mistral-7B-v0.1, and meta-llama/Llama-2-13b-hf, we used the QLoRa method, which allows converting the floating-point 32 format to smaller data types. In particular, for all three models, we used 4-bit quantization for efficient training. In the QLoRa configuration, $r$ (the rank of the low-rank matrix used in the adapters) was set to 32, $\alpha$ (the scaling factor for the learned weights) was set to 64, and dropout was set to 0.05. We applied QLoRa to all the linear layers of the models, including q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj, and lm_head. The models were trained with a batch size of 8. We used 20% of the WD1 training split as a validation set. Based on this validation set, we fixed the number of training steps to 25,000 for the pairwise models and 1,500 for the pointwise models. Note that we need fewer training steps for the pointwise model, because each mini-batch consists of 8 entities, and we consider all pairwise combinations of these entities. In contrast, for the pairwise model, each mini-batch consists of 8 pairwise combinations. We also observed that the pointwise model converges more quickly than the pairwise model.
@@
Table 5 show for how many cases ChatGPT and GPT-4 failed to answer with yes or no, when asked about pairwise comparisons. Overall, such cases were rare. The highest number of failures were seen for TG dataset, which appears to be related to the subjective nature of the features involved.
@@
Table 6 gives an overview of the properties that were selected for the WD1 and WD2 datasets, along with the corresponding number of entities. Table 7 and 8 similarly show the tags that have been considered for the Movies and Books datasets, along with the number of corresponding entities.
@@
Sweetness
coulommiers cheese, chaource cheese, mimolette cheese, pasta with soy sauce, latte without sugar, reblochon cheese, plain yogurt, mont d’or cheese, saint-agur cheese, faisselle
martini with lemon juice, martini, cod fritters, champignons crus with vinaigrette, hamburger, salty crackers, light lager, andouillette sausage, soft boiled eggs, omelette with vinegar
Saltiness
marinated mussels, liquorice candy, fortified wines, kir, oriental pastries, cola soda, petit suisse with sugar, petit suisse with sugar and cream, aperitif with anise, petit suisse
carrot puree with cream, mix vegetables salad, moussaka, guacamole, pies, zucchini, stuffed zucchini, quiches, bulgur, broccoli with cream
Scary
Pirates of the Caribbean: The Curse of the Black Pearl (2003), Interview with the Vampire: The Vampire Chronicles (1994), Scream (1996), Terminator, The (1984), Quills (2000), Batman Begins (2005), Evil Dead II (Dead by Dawn) (1987), Dawn of the Dead (1978), Spirited Away (Sen to Chihiro no kamikakushi) (2001), Requiem for a Dream (2000)
Final Fantasy: The Spirits Within (2001), Eye of the Needle (1981), Sunless (Sans Soleil) (1983), Outland (1981), Super Size Me (2004), Slumdog Millionaire (2008), Underworld (2003), Roger & Me (1989), One Hour Photo (2002), Close Encounters of the Third Kind (1977)
Funny
Fargo (1996), Original Kings of Comedy, The (2000), Meet the Spartans (2008), Simpsons Movie, The (2007), Happy Gilmore (1996), Jackass Number Two (2006), Who Framed Roger Rabbit? (1988), Tenacious D in The Pick of Destiny (2006), Men in Black (a.k.a. MIB) (1997), Elf (2003)
Ref, The (1994), American Psycho (2000), Run Lola Run (Lola rennt) (1998), Charter Trip, The (a.k.a. Package Tour, The) (1980), Bend It Like Beckham (2002), License to Drive (1988), Battleship Potemkin (1925), Jesus Camp (2006), Slap Shot (1977), Night of the Living Dead (1968)
Population
Djibouti, Qatar, Eritrea, Botswana, Papua New Guinea, Gabon, Libya, Mongolia, Mauritania, Namibia
Burundi, Rwanda, Switzerland, Wales, Kingdom of the Netherlands, Belgium, England, Italy, Netherlands, Czech Republic
Table 9: Error analysis, showing the entities with the maximum difference in rank position between the ground truth ranking and the predicted ranking.

### v2

@@ 5 Experiments
Table 4 shows the 10 highest and lowest ranked entities for some selected features, according to the rankings from the SVM method with the pairwise Llama3-8B model. The results for sweetness and saltiness were obtained with the model that was trained on WD+TG+Rocks. The rankings for scary and funny movies were obtained with the model that was trained on WD+Taste+Rocks. The ranking for population was obtained with the model that was trained on WD1. The table shows that the model was successful in selecting these top and bottom ranked entities. The top-ranked entities for sweetness, for instance, are all clearly sweet food items, while none of the bottom ranked entities are. Similar observations can be made for the other features. The model is sometimes less successful in distinguishing middle-ranked entities from bottom-ranked entities. For instance, most cheeses appear at the bottom of the ground truth ranking, whereas the model predicted these to be somewhere closer to the middle. [^11] A more detailed analysis can be found in the appendix. For population, we can see that while the top-ranked entities are all countries with a high population, their relative ranking is not accurate.
To further test the ability of the model to generalise to unseen properties, Table 5 shows results that were obtained for a number of selected features, focusing primarily on commonsense properties. For instance, for the first example, we compared a number of items according to their suitability as a present for a 10-year-old girl. To obtain the rankings, we exhaustively compared every pair of entities (from the considered set) and used the SVM method. The results were obtained with the pairwise Llama3-8B model that was trained on WD+TG+Rocks. These examples illustrate that the model often struggles with commonsense features. For instance, spaceship is listed as the most suitable vehicle for long-distance travel with a family, the school location is considered to be the least important criterion for selecting a primary school and a bathroom renovation is considered to be more urgent than fixing a leak. The disappointing results for commonsense features are somewhat surprising, although Chatterjee and Schockaert (2023) similarly found that ChatGPT struggled with commonsense properties.
@@ 6 Conclusions
Acknowledgments
This work was supported by EPSRC grants EP/V025961/1 and EP/W003309/1.
@@ Appendix A Further Details
To train the four base models, we used the QLoRa method, which allows converting the floating-point 32 format to smaller data types. In particular, for all three models, we used 4-bit quantization for efficient training. In the QLoRa configuration, $r$ (the rank of the low-rank matrix used in the adapters) was set to 32, $\alpha$ (the scaling factor for the learned weights) was set to 64, and dropout was set to 0.05. We applied QLoRa to all the linear layers of the models, including q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj, and lm_head. The models were trained with a batch size of 8. We used 20% of the WD1 training split as a validation set. Based on this validation set, we fixed the number of training steps to 12500 for the pairwise models and 1500 for the pointwise models. Note that we need fewer training steps for the pointwise model, because each mini-batch consists of 8 entities, and we consider all pairwise combinations of these entities. In contrast, for the pairwise model, each mini-batch consists of 8 pairwise combinations. We also observed that the pointwise model converges more quickly than the pairwise model.
@@
Table 6 shows for how many cases ChatGPT and GPT-4 failed to answer with yes or no, when asked about pairwise comparisons. Overall, such cases were rare. The highest number of failures was seen for TG dataset, which appears to be related to the subjective nature of the features involved.
@@
Table 7 gives an overview of the properties that were selected for the WD1 and WD2 datasets, along with the corresponding number of entities. Table 8 and 9 similarly show the tags that have been considered for the Movies and Books datasets, along with the number of corresponding entities.
Few-shot Baseline
For the few-shot configuration in Table 2, we use the pre-trained Llama3-8B model with the following prompt:
Answer the following with Yes or No only. In the worst case, if you do not know the answer then choose randomly between Yes and No.
This question is about two rivers: Is Nile longer than Indus?
Yes
This question is about two countries: Is Japan more populated than India?
No
This question is about two countries: Is China larger than India?
Yes
Note that the same three in-context examples were provided for all datasets.
@@
Sweetness coulommiers cheese, chaource cheese, mimolette cheese, pasta with soy sauce, latte without sugar, reblochon cheese, plain yogurt, mont d’or cheese, saint-agur cheese, faisselle martini with lemon juice, martini, cod fritters, champignons crus with vinaigrette, hamburger, salty crackers, light lager, andouillette sausage, soft boiled eggs, omelette with vinegar
Saltiness marinated mussels, liquorice candy, fortified wines, kir, oriental pastries, cola soda, petit suisse with sugar, petit suisse with sugar and cream, aperitif with anise, petit suisse carrot puree with cream, mix vegetables salad, moussaka, guacamole, pies, zucchini, stuffed zucchini, quiches, bulgur, broccoli with cream
Scary Pirates of the Caribbean: The Curse of the Black Pearl (2003), Interview with the Vampire: The Vampire Chronicles (1994), Scream (1996), Terminator, The (1984), Quills (2000), Batman Begins (2005), Evil Dead II (Dead by Dawn) (1987), Dawn of the Dead (1978), Spirited Away (Sen to Chihiro no kamikakushi) (2001), Requiem for a Dream (2000) Final Fantasy: The Spirits Within (2001), Eye of the Needle (1981), Sunless (Sans Soleil) (1983), Outland (1981), Super Size Me (2004), Slumdog Millionaire (2008), Underworld (2003), Roger & Me (1989), One Hour Photo (2002), Close Encounters of the Third Kind (1977)
Funny Fargo (1996), Original Kings of Comedy, The (2000), Meet the Spartans (2008), Simpsons Movie, The (2007), Happy Gilmore (1996), Jackass Number Two (2006), Who Framed Roger Rabbit? (1988), Tenacious D in The Pick of Destiny (2006), Men in Black (a.k.a. MIB) (1997), Elf (2003) Ref, The (1994), American Psycho (2000), Run Lola Run (Lola rennt) (1998), Charter Trip, The (a.k.a. Package Tour, The) (1980), Bend It Like Beckham (2002), License to Drive (1988), Battleship Potemkin (1925), Jesus Camp (2006), Slap Shot (1977), Night of the Living Dead (1968)
Population Djibouti, Qatar, Eritrea, Botswana, Papua New Guinea, Gabon, Libya, Mongolia, Mauritania, Namibia Burundi, Rwanda, Switzerland, Wales, Kingdom of the Netherlands, Belgium, England, Italy, Netherlands, Czech Republic
Table 10: Llama2-13B error analysis, showing the entities with the maximum difference in rank position between the ground truth ranking and the predicted ranking.
@@ Appendix B Additional Analysis
Wikidata Taste Rocks TG Phys
WD1-test
WD2
Sweetness
Saltiness
Sourness
Bitterness
Umaminess
Fattiness
Lightness
Grain size
Roughness
Shininess
Organisation
Variability
Density
Movies
Books
Size
Height
Mass
WD+TG+Rocks - - 74.0 72.4 60.0 60.2 70.6 72.2 - - - - - - - - - 85.6 88.2 62.0
WD+Taste+Rocks - - - - - - - - - - - - - - - 69.1 65.2 89.4 91.2 63.8
[…]
Sweetness
Saltiness
Sourness
Bitterness
Umaminess
Fattiness
Pointwise 51.0 64.8 32.5 35.2 52.0 61.7
SVM (5 samples) 62.1 62.2 42.6 44.6 56.4 63.0
SVM (30 samples) 66.0 64.7 47.6 47.0 60.6 65.8
Count (5 samples) 59.0 57.4 46.7 41.7 53.5 60.1
Count (30 samples) 64.8 64.7 49.1 47.2 59.9 64.7
Ada 17.5 8.5 12.2 16.4 22.5 10.7
Babbage 19.5 51.1 20.2 22.0 22.6 16.0
Curie 36.0 46.3 32.8 23.2 22.6 31.7
Davinci 55.0 63.2 33.3 27.2 57.0 52.0
Table 12: Comparison of ranking strategies, in terms of Spearman $\rho$%. Baseline results marked with were taken from Chatterjee et al. (2023). All other results are obtained with Llama2-13B trained on WD+TG+Rocks.
Analysis using Llama2-13B
In the main paper, we primarily used Llama3-8B for the experiments. Tables 11 and 12 show the corresponding results for Llama2-13B, focusing respectively on analysing the impact of using different training sets and on comparing different ranking strategies. Table 14 shows the 10 highest and lowest ranked entities, according to the rankings from the SVM method with the pairwise Llama2-13B model. These examples can be contrasted with the Llama3 results from Table 4.
@@
Accuracy Variability
Table 13 shows standard deviations across four runs for one of the best-performing models. As can be seen, the standard deviations are relatively small, indicating that the main differences across the models are not due to random variations.
Taste Phys
Sweetness
Saltiness
Sourness
Bitterness
Umaminess
Fattiness
Size
Height
Mass
WD+TG+Rocks 73.8 $\pm$ 0.6 81.8 $\pm$ 0.6 63.7 $\pm$ 2.5 57.9 $\pm$ 0.6 72.9 $\pm$ 0.3 70.0 $\pm$ 0.4 89.9 $\pm$ 0.9 89.2 $\pm$ 0.8 58.2 $\pm$ 1.4
Table 13: Mean accuracy and standard deviations across four runs for the pairwise Llama3-8B model.
Feature Top ranked entities Bottom ranked entities
Sweetness mango, dried date, white chocolate, peach , pineapple in syrup, fruit candy, syrup with water, ice cream, strawberry, sweet pancake with maple syrup minced beef patty, grilled calf livers, squid, sandwich with cold cuts, gizzards, croque-monsieur, roast rabbit, stir-fried bacon, roast beef , calf head with vinaigrette
Saltiness green olives, extruded salty crackers, soy sprouts with soy sauce, canned anchovies, canned sardines, pasta with soy sauce, salted pies, marinated mussels, potato chips, salted cake clafoutis, raspberry cake, stewed apple, raspberry with whipped cream, white chocolate, strawberry with cream and sugar, mix fruits juice, apple, raspberry, strawberry
Scary Descent, The (2005), Grudge, The (2004), Exorcist, The (1973), Silence of the Lambs, The (1991), Ring, The (2002), Texas Chainsaw Massacre, The (1974), Shining, The (1980), Seven (a.k.a. Se7en) (1995), Amityville Horror, The (2005), American Werewolf in London, An (1981) Super Size Me (2004), Station Agent, The (2003), Ray (2004), Dances with Wolves (1990), Jerry Maguire (1996), Driving Miss Daisy (1989), School of Rock (2003), Kung Fu Panda (2008), Miss Congeniality (2000), Ninotchka (1939)
Funny Ace Ventura: When Nature Calls (1995), Ace Ventura: Pet Detective (1994), Hot Shots! Part Deux (1993), Army of Darkness (1993), South Park: Bigger, Longer and Uncut (1999), Auntie Mame (1958), Blazing Saddles (1974), Clerks (1994), Grand Day Out with Wallace and Gromit, A (1989), Hitchhiker’s Guide to the Galaxy, The (2005) Spanish Prisoner, The (1997), Son of Dracula (1943), Ghost Dog: The Way of the Samurai (1999), Ferngully: The Last Rainforest (1992), High Crimes (2002), Cadillac Man (1990), Bad Boys II (2003), House of Wax (1953), Fire in the Sky (1993), Step Up 2 the Streets (2008)
Population India, Nigeria, People’s Republic of China, Iran, Pakistan, United States of America, Russia, Indonesia, Egypt, Bangladesh Dominica, Nauru, Andorra, Cook Islands, Saint Vincent and the Grenadines, Seychelles, Palau, Northern Mariana Islands, Liechtenstein, Niue
Table 14: We show the top and bottom ranked entities for five features: sweetness and saltiness from the food dataset, scary and funny from movies, and countries population from WD2. Results were obtained with the pairwise Llama2-13B model trained on WD+TG+Rocks for sweetness and saltiness, trained on WD+Taste+Rocks for scary and funny, and trained on WD1 for population.